OpenStudy (anonymous):

Simplify the expression 3/a^-9 b A. 27a/b B. 3/ab^-8 C. 3a^9/b

OpenStudy (jiteshmeghwal9):

\[\LARGE{{3}\over{\frac{1}{a}}^{9b}}\]

OpenStudy (anonymous):

thats not the answer. this is the problem

OpenStudy (anonymous):

thats the expression

OpenStudy (anonymous):

3 a^9 b a = 0 b = 0 integral_(-L)^L integral_(-L)^L 3 a^9 b db da = 0

OpenStudy (anonymous):

for steps visit http://www.wolframalpha.com/input/?i=3%2Fa%5E-9+b+

OpenStudy (anonymous):

yes i found the ans

OpenStudy (anonymous):

got it is c